OpenStudy (anonymous):

Solve for x, given the equation Square root of x minus 5 + 7 = 11. x = 21, solution is extraneous x = 21, solution is not extraneous x = 81, solution is extraneous x = 81, solution is not extraneous

OpenStudy (anonymous):

\[\sqrt{x-5}+7=11\]

TheSmartOne (thesmartone):

We can first subtract on both sides so that we have isolated the squareroot

OpenStudy (anonymous):

that might've been what I did wrong I started off doing 11^2 \[\sqrt{x-5}=4\]

TheSmartOne (thesmartone):

Now if you square both sides, you will result with: \(\sf (\sqrt{x-5})^2 = 4^2\\x-5= 16\)

OpenStudy (anonymous):

then add 5 to both sides thatll be x=21 now whats an extraneous solution

TheSmartOne (thesmartone):

An extraneous solution is one that will not work if you plug it back in. So we need to test if x = 21 actually works.

OpenStudy (mathmale):

Recall that you will obtain TWO possible solutions; it's up to you to determine which one is a solution and which is an extraneous solution.

OpenStudy (anonymous):

\[\sqrt{21-5}=4\]

TheSmartOne (thesmartone):

That means it is not an extraneous solution. :)

OpenStudy (anonymous):

wouldn't it be extraneous? 21-5=16 then square root is 4

TheSmartOne (thesmartone):

and then 4 + 7 = 11

TheSmartOne (thesmartone):

http://hotmath.com/hotmath_help/topics/extraneous-solutions.html "An extraneous solution is a root of a transformed equation that is not a root of the original equation because it was excluded from the domain of the original equation."

OpenStudy (anonymous):

So it's not?

TheSmartOne (thesmartone):

Yes, it is not extraneous.

OpenStudy (anonymous):

okay thank you for your help

TheSmartOne (thesmartone):

Anytime! :)